OpenStudy (anonymous):

The sum of the digits in a two-digit number is 7. The new number obtained when the digits are reversed is 27 less than the original number. Find the original number.

OpenStudy (diyadiya):

Let the two digit number be xy Its given that the sum of the digits in a two-digit number is 7 So what do you think will be the first equation ?

OpenStudy (anonymous):

x + y = 7 I had that part.....I just go wrong from there. Does is go something like (10x + y) - (10y + x) ?

OpenStudy (diyadiya):

Yeah you are almost right But check the question again ,its given "The new number obtained when the digits are reversed is 27 less than the original number." original number is 10x+y and number obtained after reversing is 10y+x Now what will be the second equation ? **we are multiplying by 10 since the digit is in the ten's place

OpenStudy (anonymous):

So it would be like 10x + y (original) - 27 = 10y + x?

OpenStudy (diyadiya):

Right!! :D Now i think you can reduce the equation to a more simpler form? can you?

OpenStudy (anonymous):

Like this? 10x + y - 27 = 10y + x -------> 9x-27=9y Could 9x - 27 become 9(x-3)? divide by 9 and y=x-3?

OpenStudy (diyadiya):

Exactly !!:D

OpenStudy (anonymous):

then plug it in to our first equation? or second.....

OpenStudy (anonymous):

It's 52!!! We did it!

OpenStudy (diyadiya):

Yay !!! Good!!! :D

OpenStudy (anonymous):

Can you help me with some others?

OpenStudy (diyadiya):

Sure :D